### v1

@@ 2.3 Quantifying the Sampling Spread
Figure 5: Sampling spread and generalization to unseen climates of the generative probabilistic downscaling process. (A) The ESM field interpolated to the target resolution. (B) and (C) show two different exemplary samples generated by the CM downscaling, preserving large-scale patterns and generating new patterns on smaller scales. (D) The ensemble mean of $10^{3}$ samples with the standard deviation shown in (E). (F) Three-year rolling global mean normalized to the reference year 2020 of the extreme warming scenario SSP5-8.5. The ESM (orange) shows an increase in global mean precipitation over the century in line with the thermodynamic Clausius-Clapeyron relation [Traxl et al., 2021]. The CM downscaling (magenta) is able to preserve the trend with a high degree of accuracy, notably without the addition of any physical constraints in the CM network.
@@ 2.4 Future Projections
The efficient nature of our CM approach enables the downscaling of long climate simulations over a century, such as prescribed with the SPP5-8.5 extreme warming scenario from the Coupled Model Intercomparison Project Phase 6 (CMIP6). We find that the downscaling accurately preserves the global precipitation content in the ESM and preserves the non-linear trends of the warming scenario (Fig. 5F). The network thus generalizes to out-of-sample predictions without the need for hard-coding auxiliary physical constraints in the network as, for example, done in [Beucler et al., 2021, Harder et al., 2022, Hess et al., 2022].
@@ 3 Discussion
We introduced a generative foundation model for efficient, scale-adaptive, and probabilistic downscaling of ESM precipitation fields. Our approach is based on a novel generative machine learning method that learns a self-consistent approximation of a reversed diffusion process. The CM is able to generate highly realistic global precipitation fields learned from the ERA5 reanalysis, which is employed as ground truth, in a single step. Our framework corrects the representation of extreme events as well as spatial ESM patterns especially at the small spatial scales, which are both crucial for impact assessments. Moreover, spatial biases are also corrected efficiently.
@@
The CM method is up to three orders of magnitude faster than current diffusion models based on SDEs, which need to solve a differential equation iteratively. Crucially, the CM maintains a high degree of controllability to guide the sampling in such a way that spatial patterns larger than a chosen spatial scale in Earth system models are preserved.
@@
Similar to the SDE-based method that [Bischoff and Deck, 2023] apply to idealized fluid dynamics, our CM-based foundation model only needs to be trained once on a given high-resolution observational target dataset. Since we do not condition on ESM input during training, our method can be applied to any ESM without the need for retraining. Combined with the efficient sampling of CMs, our approach is computationally cheap and fast, particularly when processing large ensembles of ESMs, without noticible trade-offs in accuracy.
@@
Our CM-based approach can create highly realistic fields that maintain high correlation levels with the ESM. We find comparable or better results when compared with the much more computationally expensive SDE-based method. The efficient single-step generation of ESM fields will be particularly relevant for processing large datasets, e.g., large ensembles as needed for uncertainty quantification, for simulations with a high temporal resolution, or long-term studies such as those in paleo-climate simulations. Improving the efficiency of current deep learning models is also important from an energy consumption perspective; in this regard our method provides a valuable contribution towards “greener AI” [Schwartz et al., 2019].
@@
When evaluated on a future extreme emission scenario, we find that our generative CM method accurately preserves the trend of the ESM, even for an extreme scenario of greenhouse gas emissions and associated global warming. This is remarkable since many machine learning-based applications to climate dynamics struggle with the out-of-sample problem imposed by our highly non-stationary climate system, when trained on historical data alone. In contrast to previous studies [Beucler et al., 2021, Harder et al., 2022, Hess et al., 2022], to achieve this it is not necessary to add specifically formulated physical constraints to our model. Our unconstrained CM method hence allows for a more natural generalization to unseen climate states, inherently translating non-stationary dynamics from the ESM to the downscaled high-resolution fields.
@@
We showcase our method on uni-variate precipitation simulations, because precipitation is arguably the most difficult to model climate variable. An extension to multi-variate downscaling is a natural extension of our study in future research. In principle, the convolutional CM network can be extended to include further variables as additional channels in a straightforward manner. The consistency scale will then depend on the variable (channel) and might require a separate treatment, which we leave for future research.
Acknowledgments
The authors thank Katherine Deck for her valuable input. NB and PH acknowledge funding by the Volkswagen Foundation, as well as the European Regional Development Fund (ERDF), the German Federal Ministry of Education and Research, and the Land Brandenburg for supporting this project by providing resources on the high-performance computer system at the Potsdam Institute for Climate Impact Research. MA acknowledges funding under the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare. This is ClimTip contribution #X; the ClimTip project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101137601.
Data Availability
The ERA5 reanalysis data is available for download at the Copernicus Climate Change Service (C3S) (https://cds.climate.copernicus.eu/cdsapp#!/dataset/reanalysis-era5-single-levels?tab=overview). The simulation data from the POEM ESM is available at https://doi.org/10.5281/zenodo.4683086.
Code Availability
The Python code for processing and analyzing the data, together with the PyTorch [Paszke et al., 2019] code for training, will be made available on GitHub: https://github.com/p-hss/consistency-climate-downscaling.git.
Competing interests
The authors declare no competing interests.
@@ 4.1 Training Data
As a target and ground truth dataset, we use observational precipitation data from the ERA5 reanalysis [Hersbach et al., 2020] provided by the European Center for Medium-Range Weather Forecasting (ECMWF). We bilinearly interpolate the reanalysis data to a resolution of $0.75^{\circ}$ and $0.9375^{\circ}$ in latitude and longitude direction, respectively (i.e. $240\times 384$ grid points), which corresponds to a four times higher resolution compared to the raw Earth system model simulations with $3^{\circ}\times 3.75^{\circ}$ resolution (i.e. 60×96)60$\times$ 96)60 × 96 ) grid points).
For the ESM precipitation fields, we use global simulations from the fully coupled Potsdam Earth Model (POEM) [Drüke et al., 2021], which includes model components for the atmosphere, ocean, ice sheets, and dynamic vegetation. Both ESM and reanalysis datasets cover the period from 1940 to 2018, and we split the data into a training set from 1940-1990, a validation set from 1991-2003, and a test set from 2004-2018.
We apply several preprocessing steps to the ESM data. We interpolate the POEM simulations onto the same high-resolution grid as the ground truth ERA5 data for downscaling purposes and model evaluation. A low-pass filter is then applied to remove small-scale artifacts created by the interpolation. Quantile delta mapping (QDM) [Cannon et al., 2015] with 500 quantiles is applied in a standard way to remove distributional biases in the ESM simulation for each grid cell individually. As discussed in section 2.2, the generative downscaling only corrects biases related to a specified spatial scale. Hence, the QDM step ensures a strong reduction of single-cell biases, while the generative downscaling corrects spatial patterns that are physically consistent. Finally, the ESM and ERA5 data are log-transformed, $\tilde{x}=\log(x+\epsilon)-\log(\epsilon)$ with $\epsilon=0.0001$, followed by a normalization approximately into the range [-1, 1].
@@ 4.2 Score-based Diffusion Models
The underlying idea of diffusion-based generative models is to learn a reverse diffusion process from a known prior distribution $\mathbf{x}(t=T)\sim p_{T}$, such as a Gaussian distribution, to a target data distribution $\mathbf{x}(t=0)\sim p_{0}$, where $\mathbf{x}\in\mathbb{R}^{d}$ and $d$ is the data dimension, e.g., the number of pixels in an image. Score-based generative diffusion models [Song and Ermon, 2019, Song et al., 2021, Song et al., 2022] generalize probabilistic denoising diffusion models [Ho et al., 2020, Dhariwal and Nichol, 2021] to continuous-time stochastic differential equations (SDEs).
@@
where $\mu(\mathbf{x},t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift term, $\mathbf{w}$ denotes a Wiener process and $g(t):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient. The reverse SDE used to generate images from noise is given by [Song et al., 2021]
@@
with $\bar{t}$ denoting a time reversal and $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ being the score function of the target distribution. The score function is not analytically tractable, but one can train a score network, $s(\mathbf{x},t;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to approximate the score function $s(\mathbf{x},t;\bm{\phi})\approx\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$, e.g., using denoising score matching [Song et al., 2022] (see SI for details). For sampling, we use the Euler-Maruyama solver to integrate the reverse SDE from $t=T$ to $t=0$ in Eq. 2 with $500$ steps.
@@ 4.3 Consistency Models
One major drawback of current diffusion models is that the numerical integration of the differential equation requires around 10-2000 network evaluations, depending on the solver. This makes the generation process computationally inefficient and costly compared to other generative models such as GANs [Goodfellow et al., 2014] or NFs [Dinh et al., 2015, Papamakarios et al., 2021], which can generate images in a single network evaluation. Distillation techniques can reduce the number of integration steps of diffusion models, which often represent a computational bottleneck [Luhman and Luhman, 2021, Zheng et al., 2023].
Consistency models (CMs) can be trained from scratch without distillation and only require a single step to generate a new sample. They have been shown to outperform current distillation techniques [Song et al., 2023]. CMs learn a consistency function, $f(\mathbf{x}(t),t)=\mathbf{x}(t_{\mathrm{min}})$, which is self-consistent, i.e.,
@@
where the time interval is here set to $t_{\mathrm{min}}=0.002$ and $t_{\mathrm{max}}=80$. Further, a boundary condition $f(\mathbf{x}(t_{\mathrm{min}}),t_{\mathrm{min}})=\mathbf{x}(t_{\mathrm{min}})$, for $t=t_{\mathrm{min}}$ is imposed. This can be implemented with the parameterization:
@@
where the coefficients are defined, following [Karras et al., 2022, Song et al., 2023], as
@@
where $\mathbb{E}_{\mathbf{x},n,\mathbf{x}_{t_{n}}}\equiv\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}},t_{n}\sim\mathcal{U}(1,N(k)-1),\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{1})}$, and $N(\cdot)$ increases over training time with a given schedule. With $\bm{\bar{\theta}}$ we denote an exponential moving average (EMA) over the model parameters $\bm{\theta}$, updated with $\bm{\bar{\theta}}=\mathrm{stopgrad}[w(k)\bm{\bar{\theta}}+(1-w(k)\bm{\theta})]$, and a schedule $w(\cdot)$ for the decay rate.
For the distance measure $d(\cdot,\cdot)$, we follow [Song et al., 2023] and use a combination of the learned perceptual image patch similarity (LPIPS) [Zhang et al., 2018] and $l^{1}$ norm:

### v3

@@ 2.3 Quantifying the Sampling Spread
We evaluate an ensemble of 100 downscaled realizations of coarse ERA5 fields using the continuous ranked probability score (CRPS) [hersbach_decomposition_2000], for three different noise scales, corresponding to $t=t_{\text{min}}$, $t=t^{*}$, and $t=t_{\text{max}}$ (Fig. S7). Further, the CRPS is evaluated as a skill score (CRPSS) relative to the baseline ensemble of 100 random high-resolution ERA5 fields. We find that the noise strength enables calibration of the ensemble spread, which is sharp for small noise scales and increasing broad for larger noise scales. The intermediate noise scale $t=t^{*}$ shows the lowest CRPS and maximum noise leads to a similar CRPSS to the baseline climatology of drawing random fields. For definitions of the CRPS and CRPSS see section LABEL:ext-sec:ens_spread in the SI.
Figure 5: Sampling spread and generalization to unseen climates of the generative probabilistic downscaling process. (A) The ESM field interpolated to the target resolution. (B) and (C) show two different exemplary samples generated by the CM downscaling, preserving large-scale patterns and generating new patterns on smaller scales. (D) The ensemble mean of $10^{3}$ samples with the standard deviation is shown in (E). (F) Three-year rolling global mean normalized to the reference year 2020 of the very high emission scenario SSP5-8.5. The ESM (orange) shows an increase in global mean precipitation over the emission scenario, in line with the thermodynamic Clausius-Clapeyron relation [traxl_role_2021]. The CM downscaling (magenta) is able to preserve the trend with a high degree of accuracy, notably without the addition of any physical constraints in the CM network.
@@ 2.4 Future Projections
The efficient nature of our CM approach enables the downscaling of long climate simulations over a century, such as prescribed with the SPP5-8.5 high-emission scenario from the Coupled Model Intercomparison Project Phase 6 (CMIP6). We find that the downscaling accurately preserves the global precipitation content of the ESM and preserves the non-linear trends, as expected from the Clausius-Clapeyron relation [traxl_role_2021] (Fig. 5F). The network thus generalizes to the out-of-sample predictions of increasing global mean precipitation without the need for hard-coding auxiliary physical constraints in the network as, for example, done in [beucler_enforcing_2021, harder_hard-constrained_2023, hess_physically_2022]. However, our method can, in principle, be extended to such constraints to preserve trends exactly. We note that the ability to preserve trends is expected to be related to the noise variance used for conditioning the downscaling, as maximum noise corresponds to sampling from the learned historical ERA5 distribution, and no noise does not alter the ESM simulation in any way.
@@ 3 Discussion
We introduced a generative machine learning approach for efficient, scale-adaptive, and probabilistic downscaling of ESM precipitation fields. Our approach is based on consistency models (CMs), a recently developed method that learns a self-consistent approximation of a reversed diffusion process. The CM is able to generate highly realistic global precipitation fields learned from the ERA5 reanalysis, which is employed as ground truth, in a single step. Our framework corrects the representation of extreme events as well as spatial ESM patterns especially at the small spatial scales, which are both crucial for impact assessments. Moreover, spatial biases are also corrected efficiently.
@@
For the specific task considered here, the CM method is up to three orders of magnitude faster than current diffusion models based on SDEs [bischoff_unpaired_2024, dhariwal_diffusion_2021, song_score-based_2021], which need to solve a differential equation iteratively. Crucially, the CM maintains a high degree of controllability to guide the sampling in such a way that unbiased spatial patterns larger than a chosen spatial scale in ESMs are preserved.
@@
Similar to the SDE-based method that [bischoff_unpaired_2024] apply to idealized fluid dynamics, our CM-based model only needs to be trained once on a given high-resolution observational target dataset. Since we do not condition on ESM input during training, our method can be applied to any ESM without the need for retraining. Combined with the efficient sampling of CMs, our approach is computationally fast, particularly when processing large ensembles of ESMs, without noticeable trade-offs in accuracy.
@@
Our CM-based approach can create highly realistic fields that maintain high correlation levels with the ESMs at the spatial scales chose to be preserved. We find slightly better and at least competitive results when compared with the much more computationally expensive SDE-based method. The efficient single-step generation of ESM fields will be particularly relevant for processing large datasets, e.g., large ensembles as needed for uncertainty quantification, for simulations with a high temporal resolution, or long-term studies such as those in paleo-climate simulations. The ability to calibrate the ensemble sharpness and spread using different noise scales could be useful for weather prediction tasks as well. When applying our method to weather predictions, changes in the training data and splits might be of advantage to avoid distributional shifts due to the availability of satellite data. Improving the efficiency of current deep learning models is also important from an energy consumption perspective; in this regard, our method provides a valuable contribution towards “greener AI” [schwartz_green_2019].
@@
When evaluated on a future very high emission scenario, we find that our generative CM method accurately preserves the trend of the ESM, even for an extreme scenario of greenhouse gas emissions and associated global warming. This is remarkable since many machine learning-based applications to climate dynamics struggle with the out-of-sample problem imposed by our highly non-stationary climate system, when trained on historical data alone. In contrast to previous studies [beucler_enforcing_2021, hess_physically_2022], to achieve this it is not necessary to add specifically formulated physical constraints to our model. Our unconstrained CM method hence allows for a more natural generalization to unseen climate states, inherently translating non-stationary dynamics from the ESM to the downscaled high-resolution fields. However, since the global mean of the ESM is only approximately conserved with our method, adding a hard architecture constrained as in [harder_hard-constrained_2023], which fulfills a constraint up to machine precision, could further improve our results.
Grid-cell-wise autocorrelation estimates between subsequent fields suggest that CM-based stochastic downscaling method captures temporal variability of the high-resolution target data although not explicitly using temporal information. Extending the approach to temporal conditioning or predicting several time steps as done, e.g., in [price_gencast_2024], or using correlated noise as in [huang_blue_2024], might further improve the temporal dynamics.
Our downscaling method increases the resolution of the ESM simulations by a factor of four in this study. Given a higher resolution training target and more computational resources, it should, in principle, also allow for larger downscaling ratios.
@@
We showcase our method on uni-variate precipitation simulations because precipitation is arguably the most difficult climate variable to model. An extension to multi-variate downscaling is a natural extension of our study in future research. In principle, the convolutional CM network can be extended to include further variables as additional channels in a straightforward manner. The consistency scale will then depend on the variable (channel) and might require a separate treatment. Similarly, we believe that our method can be applied on other time scales, e.g. hourly or monthly as well, using suitably adjusted noise scales. Choosing the optimal noise scale using the intersection of power spectral densities as done here and in [bischoff_unpaired_2024] assumes that the spectral densities are monotonically decreasing with the wavenumber. While we believe that this should hold for most climate impact variables, there might be exceptions that would require a different approach. We leave these explorations for future research.
@@ 4.1 Training Data
As a target and ground truth dataset, we use observational precipitation data from the ERA5 reanalysis [hersbach_era5_2020] provided by the European Center for Medium-Range Weather Forecasting (ECMWF). It covers the period from 1940 to the present, and we split the data into a training set from 1940-1990, a validation set from 1991-2003, and a test set from 2004-2018. We bilinearly interpolate the reanalysis data to a resolution of $0.75^{\circ}$ and $0.9375^{\circ}$ in latitude and longitude direction, respectively (i.e. $240\times 384$ grid points), which corresponds to a four times higher resolution compared to the raw ESM simulations with $3^{\circ}\times 3.75^{\circ}$ resolution (i.e. 60×96)60$\times$ 96)60 × 96 ) grid points).
For the ESM precipitation fields, we use global simulations from three different ESM with varying complexity and resolution. The fully coupled Potsdam Earth Model (POEM) [druke_cm2mc-lpjml_2021], which includes model components for the atmosphere, ocean, ice sheets, and dynamic vegetation, is used as the primary model for comparison of the two generative downscaling methods for past and future climates. To demonstrate the ability of our CM-based method to correct any ESM with a coarser native resolution than the training ground truth, we further include daily precipitation simulations from the much more comprehensive and complex GFDL-ESM4 [dunne_gfdl_2020], with a native resolution of $1^{\circ}\times 1^{\circ}$. We initially upscale the GFDL-ESM4 resolution to the same grid as the POEM ESM to allow direct comparisons. We further include SpeedyWeather.jl [klower_speedyweatherjl_2024], with a native resolution of $3.75^{\circ}\times 3.75^{\circ}$, which only has a dynamic atmosphere and is, hence, less comprehensive than the fully coupled POEM ESM. Finally, we also use ERA5 data upscaled to the native POEM resolution as test data for which a paired ground truth is available. Applying our method to the latter can hence be seen as a proof of concept.
For evaluation, we use 14 years of available historical data from each of the simulations, with periods 2004-2018 for POEM, 2000-2014 for GFDL-ESM4, 1956-1970 for SpeedyWeather.jl, and 2007-2021 for ERA5.
We apply several preprocessing steps to the simulated input data. We first interpolate the input simulations onto the same high-resolution grid as the ground truth ERA5 data for downscaling purposes and model evaluation. A low-pass filter is then applied to remove small-scale artifacts created by the interpolation. Quantile delta mapping (QDM) [cannon_bias_2015] with 500 quantiles is then applied in a standard way to remove distributional biases in the ESM simulation for each grid cell individually. As discussed in section 2.2, the generative downscaling only corrects biases related to a specified spatial scale. Hence, the QDM step ensures a strong reduction of single-cell biases, while the generative downscaling corrects spatial patterns that are physically consistent. Finally, the ESMs and ERA5 data are log-transformed, $\tilde{x}=\log(x+\epsilon)-\log(\epsilon)$ with $\epsilon=0.0001$, as in [hess_physically_2022], followed by a normalization approximately into the range [-1, 1].
@@ 4.2 Score-based Diffusion Models
The underlying idea of diffusion-based generative models is to learn a reverse diffusion process from a known prior distribution $\mathbf{x}(t=T)\sim p_{T}$, such as a Gaussian distribution, to a target data distribution $\mathbf{x}(t=0)\sim p_{0}$, where $\mathbf{x}\in\mathbb{R}^{d}$ and $d$ is the data dimension, e.g., the number of pixels in an image. Score-based generative diffusion models [song_generative_2019, song_score-based_2021, song_denoising_2022] generalize probabilistic denoising diffusion models [ho_denoising_2020, dhariwal_diffusion_2021] to continuous-time stochastic differential equations (SDEs).
@@
where $\mu(\mathbf{x},t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift term, $\mathbf{w}$ denotes a Wiener process and $g(t):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient. The reverse SDE used to generate images from noise is given by [song_score-based_2021]
@@
with $\bar{t}$ denoting a time reversal and $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ being the score function of the target distribution. The score function is not analytically tractable, but one can train a score network, $s(\mathbf{x},t;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to approximate the score function $s(\mathbf{x},t;\bm{\phi})\approx\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$, e.g., using denoising score matching [song_denoising_2022] (see SI for details). For sampling, we use the Euler-Maruyama solver to integrate the reverse SDE from $t=T$ to $t=0$ in Eq. 2 with $500$ steps.
@@ 4.3 Consistency Models
One major drawback of current diffusion models is that the numerical integration of the differential equation requires around 10-2000 network evaluations, depending on the solver. This makes the generation process computationally inefficient and costly compared to other generative models such as generative adversarial networks (GANs) [goodfellow_generative_2014] or normalizing flows (NFs) [dinh_nice_2015, papamakarios_normalizing_2021], which can generate images in a single network evaluation. Distillation techniques can reduce the number of integration steps of diffusion models, which often represent a computational bottleneck [luhman_knowledge_2021, zheng_fast_2024].
Consistency models (CMs) can be trained from scratch without distillation and only require a single step to generate a new sample. They have been shown to outperform current distillation techniques [song_consistency_2023]. CMs learn a consistency function, $f(\mathbf{x}(t),t)=\mathbf{x}(t_{\mathrm{min}})$, which is self-consistent, i.e.,
@@
where the time interval is here set to $t_{\mathrm{min}}=0.002$ and $t_{\mathrm{max}}=80$, following [song_consistency_2023]. Further, a boundary condition $f(\mathbf{x}(t_{\mathrm{min}}),t_{\mathrm{min}})=\mathbf{x}(t_{\mathrm{min}})$, for $t=t_{\mathrm{min}}$ is imposed. This can be implemented with the parameterization:
@@
where $F(\cdot)$ is a UNet with parameters $\bm{\theta}$. The time information is transformed using a sin-cosine positional embedding in the network. The coefficients $c_{\mathrm{skip}}(t)$ and $c_{\mathrm{out}}$ are defined, following [karras_elucidating_2022, song_consistency_2023], as
@@
where $\mathbb{E}_{\mathbf{x},n,\mathbf{x}_{t_{n}}}\equiv\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}},n\sim\mathcal{U}(1,N(k)-1),\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{1})}$. The discrete time step is determined via
$t_{n}=\left(t_{\text{min}}^{1/\rho}+(n-1)/(N-1)(t_{\text{max}}^{1/\rho}-t_{\text{min}}^{1/\rho})\right)^{\rho},$ (7)
where $\rho=7$ and the discretization schedule is given by
$N(k)=\left(\sqrt{k/K((s_{1}+1)^{2}-s^{2}_{0})+s^{2}_{0}}-1\right)+1,$ (8)
where $k$ is the current training step, and $K$ is the estimated total number of training steps obtained from the PyTorch Lightning library. The initial discretization steps are set to $s_{0}=2$, and the maximum number of steps to $s_{1}=150$, following [song_consistency_2023]. With $\bm{\bar{\theta}}$ we denote an exponential moving average (EMA) over the model parameters $\bm{\theta}$, updated with $\bm{\bar{\theta}}=\mathrm{stopgrad}[w(k)\bm{\bar{\theta}}+(1-w(k)\bm{\theta})]$, with the decay schedule given by
$w(k)=\exp\left(\frac{s_{0}\log w_{0}}{N(k)}\right),$ (9)
where $w_{0}=0.9$ is the initial decay rate, following [song_consistency_2023]. For the distance measure $d(\cdot,\cdot)$, we follow [song_consistency_2023] and use a combination of the learned perceptual image patch similarity (LPIPS) [zhang_unreasonable_2018] and $l^{1}$ norm:
@@
Thus, the training of the CM is self-supervised and closely related to representation learning [lessig_atmorep_2023], where a so-called "online network" $f(\cdot:\bm{\theta})$ is trained to predict the same image representation as a "target network" $f(\cdot:\bar{\bm{\theta}})$ [grill_bootstrap_2020]. Importantly, the CM is thereby not trained explicitly for the downscaling tasks, which are purely performed at the inference stage.